### v1

# EpidemIQs: Prompt-to-Paper LLM Agents for Epidemic Modeling and Analysis

Abstract
Large Language Models (LLMs) offer new opportunities to automate complex interdisciplinary research domains. Epidemic modeling, characterized by its complexity and reliance on network science, dynamical systems, epidemiology, and stochastic simulations, represents a prime candidate for leveraging LLM-driven automation. We introduce EpidemIQs, a novel multi-agent LLM framework that integrates user inputs and autonomously conducts literature review, analytical derivation, network modeling, mechanistic modeling, stochastic simulations, data visualization and analysis, and finally documentation of findings in a structured manuscript. We introduced two types of agents: a scientist agent for planning, coordination, reflection, and generation of final results, and a task-expert agent to focus exclusively on one specific duty serving as a tool to the scientist agent. The framework consistently generated complete reports in scientific article format. Specifically, using GPT 4.1 and GPT 4.1 mini as backbone LLMs for scientist and task-expert agents, respectively, the autonomous process completed with average total token usage 870K at a cost of about \$1.57 per study, achieving a 100% completion success rate through our experiments. We evaluate EpidemIQs across different epidemic scenarios, measuring computational cost, completion success rate, and AI and human expert reviews of generated reports. We compare EpidemIQs to the single-agent LLM, which has the same system prompts and tools, iteratively planning, invoking tools, and revising outputs until task completion. The comparison shows consistently higher performance of the proposed framework across five different scenarios. EpidemIQs represents a step forward in accelerating scientific research by significantly reducing costs and turnaround time of discovery processes, and enhancing accessibility to advanced modeling tools.
I Introduction
Scientists’ Efforts to automate research date back to at least the early 1970s [1, 2, 3], aiming to overcome limits on the number of ideas scientists can pursue, constraints that force them to prioritize ideas with higher predicted impact. For example, the Automated Mathematician was introduced to carry out simple mathematics research guided by large sets of heuristic rules [4, 5], and DENDRAL was proposed to assist chemists with data interpretation problems [6].
More recently, advances in autoregressive large language models (LLMs) [7, 8, 9] have demonstrated strong capabilities in question answering, problem-solving, and coding, with promising applications across science and engineering. However, LLMs still face limitations when applied to real-world tasks [10]. These constraints led to the development of LLM agents, which extend LLMs with abilities such as external tool use [11, 12, 13], chain-of-thought prompting [14], and iterative self-improvement [15, 16]. LLM agents have since been applied to diverse domains, including software engineering [17, 18], cybersecurity [19], medical diagnosis [20], chemistry [21, 22, 23], material science [24, 25, 26], computational biology [27], algorithm design [28], and healthcare [29]. Multi-agent systems have also emerged to simulate human-like research environments across disciplines. Examples include Agent Laboratory [30], Virtual Lab [31], ChemCrow [32], ResearchAgent [33], and The AI Scientist [34]. For broader surveys of LLM agent applications, see [35, 36, 37, 38].
Despite the stunning advances of LLM agents across various domains, epidemiology-related areas have not yet fully benefited from these developments [39], largely because they are among the most sophisticated and interdisciplinary fields. Network-based epidemic modeling is among the sophisticated approaches to studying the spread of infectious diseases by incorporating the structure of contact networks, which represent interactions among individuals or groups in a population [40].
Unlike traditional compartmental models, which assume homogeneous mixing, network-based models account for heterogeneous contact patterns by representing individuals as nodes and their interactions as edges (possibly weighted or time-varying) in a graph. This approach captures the realistic social and spatial structures influencing disease transmission, such as clustering, degree distribution, and community structures. By integrating network topology, these models provide insights into how connectivity patterns affect epidemic dynamics, enabling more accurate predictions of disease spread, the identification of critical transmission pathways, and the evaluation of targeted intervention strategies, such as vaccination or social distancing, in complex populations [41].
Addressing these models requires an interdisciplinary expertise: (i) stochastic-process theory to analyze the master equations and their approximations; (ii) network science to characterize and parametrize contact structures; (iii) epidemiology to constrain pathogen-specific parameters; and (iv) computational epidemiology to conduct simulations and scenario analysis. Having a research group with this level of diversity is not only challenging to access but also poses coordination challenges. In this study, we introduce EpidemIQs, a novel multi-LLM agents capable of autonomously addressing research in network-based epidemic modeling for the first time.
EpidemIQs emulates an interdisciplinary research environment with five collaborative teams of task-expert and scientist agents, capable of performing end-to-end autonomous research: given a user query, if identified as a network-based problem, it initiates the process by discovering relevant insights from diverse online sources (e.g., web-based content and published papers) to conduct a literature review and address analytical aspects through mathematical reasoning. It then formulates the problem as a network-based model, tests it through rigorous stochastic simulations, analyzes the resulting multimodal data (e.g., visual and numerical), and finally composes a scientifically formatted manuscript that reports the entire procedure, from problem formulation to simulation results, discussion, and final conclusion. The framework can also operate in copilot mode, which enables human cooperation throughout the process.
We evaluate our proposed framework across five epidemic scenarios articulated through questions, three of which were unknown to the framework. We tested EpidemIQs against the multi-turn single-agent baseline, comparing performance by (i) discussing the results of each team across all scenarios, (ii) AI and human expert evaluations of generated papers, and (iii) computational statistics. The proposed framework consistently showed impressive performance, achieving a 100% completion success rate, an average human review score of 7.98 out of 10, a low cost of only \$1.57, and a total processing time of 1,190 seconds. We can summarize the main contributions of our work as follows:
  • Developed novel multi-LLM agents that conduct epidemic research through five distinct scientific phases.
  • Integrated multimodal data (visual, textual, numerical, graph structures, etc.) from diverse sources, including online retrieval, experimental results, and internally generated artifacts (e.g., graphs, tables), to enhance utility and adaptability across various research contexts.
  • Enabled fully autonomous mode to minimize human intervention, requiring only an initial query, while also supporting a copilot mode for collaborative operation.
  • Ensured high interpretability by having each scientist provide reasoning for its actions and generated outcomes, and recording results and scripts to facilitate validation of procedures and outcomes.
The remainder of this paper is organized as follows: Section II describes the methodology. Section III details the design and implementation of the system. Section IV outlines the experimental setup and evaluation cases. Section V presents the results and discusses the study’s limitations and shortcomings. Finally, Section VII concludes the paper and proposes directions for future work.
II Methodology
LLMs demonstrate advanced capabilities relevant to scientific tasks, including coding and answering technical questions. However, their performance in tackling complex epidemiology scenarios remains limited. To address these challenges, we propose a multi-agent framework that emulates a scientific epidemic research laboratory environment, designed to enhance the practical capabilities of LLMs in complex scientific tasks. The EpidemIQs framework is organized into four core functional layers that collectively support autonomous research workflows[38]: (1) The multi-agent orchestration layer governs the coordination, communication, and cooperation among specialized agents.(2) The backbone LLM operates as the central reasoning and decision-making unit . (3) The perception layer collects, integrates, and processes heterogeneous data sources, including scientific literature, experimental results, contextual information, and prior knowledge bases. (4) The action layer executes tasks derived from the LLM’s output, encompassing activities such as code generation, simulation, modeling, etc.
Figure 1: Multimodal approach for the multi-agent framework for epidemic autonomous research. This framework enhances the power of multi-LLM agents by integrating multimodal data from various sources, such as literature, online web data, theoretical models, and simulation and experiments.
Looking at Figure 1, the architectural design illustrates the structured interaction among the four functional layers that enable autonomous epidemic research. The orchestration layer governs inter-agent communication, task allocation, and coordination protocols, ensuring efficient collaboration among specialized agents. Within the perception layer, agents process multimodal data streams through standardized interfaces that convert heterogeneous sources, including scientific literature, experimental data, contextual metadata, and structured knowledge bases, into unified semantic representations suitable for downstream reasoning and analysis. The backbone LLM operates as the central cognitive engine, analyzing perceived inputs, generating hypotheses, formulating research plans, and executing decisions based on accumulated evidence and domain constraints. The agents’ outputs inform the action layer, where the system autonomously selects and calls functions to deploy tools, such as stochastic simulation engines, Retrieval-Augmented Generation (RAG), API calls, and executes it in the correct order, so that high‑level research goals are broken down into concrete, reproducible steps such as loading data, running simulations, and storing results.
III Multi-LLM Agent Architecture: EpidemIQs Framework
The orchestration of EpidemIQs is illustrated in Figure 2. This framework is structured to emulate a scientific laboratory, comprising five phases: Discovery, Modeling, Simulation, Analysis, and Report Writing. These phases are executed by distinct collaborative multi-agent system teams, and the results of their work are integrated to produce the final report. In this section, we first describe the design of two types of agents, followed by the architecture design, output structure, memory, and implementation details of EpidemIQs.
Figure 2: EpidemIQs orchestration across five research phases. Specialized agent teams perform tool calling (via red arrows), collaborate with each other (curved blue arrows), and exchange artifacts through a shared project repository (hexagons), enabling an end-to-end pipeline from discovery to final report generation.
III-A Two Agent Types: Scientists and Experts
Two agent types are inspired by the generative agent concept [42] and further advanced by frameworks such as CAMEL [43], AutoGen [44], and Metagpt [45], and are designed to enable autonomous scientific reasoning and investigation. The central coordinating component, referred to as the scientist agent, performs high-level orchestration through iterative planning, reflection, and execution refinement. Upon receiving an input query, the scientist employs a dedicated plan module, which parses and decomposes the query into a plan comprising sub-tasks and their logical dependencies. Both reflect and ReAct modules are constrained to generate output in predefined formats. While structured outputs, e.g., JSON, enhance reproducibility and reliability of performance, they have recently been shown to compromise the reasoning capability of LLMs [46]. Therefore, we explicitly devote a separate plan module as an agent to allow it to freely generate the plan, while both the subsequent ReAct and reflect modules are restricted to generating predefined structured outputs.
Figure 3: The scientist agent architecture. a) Three modules as plan, ReAct, and reflect. The reflection ends either when the reflect agent decides the answer is already good enough or it reaches its maximum iteration limit. b) the conversation components for ReAct agent, as can be seen each time with the fixed task and plan, it updates new actions, results, and reflection. At each iteration $t$, the newly generated $\text{Action}_{t}$, corresponding $\text{Result}_{t}$, and $\text{Reflection}_{t}$ are appended to short-term memory, while the $\text{Action}_{t-2}$, corresponding $\text{Result}_{t-2}$ are dropped out. The loop continues updating the conversation trace until the aforementioned conditions are met.
Once the plan is formulated, the scientist agent iteratively executes the ReAct-Reflect loop until either the reflect agent decides no further revision is needed, or it reaches the maximum iteration limit, $r_{max}$. The ReAct Module executes reasoning, and call tools conditioned on the current plan segment, and refine its response based on the action outcomes, while the reflect module prompted to perform logical reasoning, evaluates intermediate outputs for errors, logical inconsistencies, or incomplete results, producing structured JSON feedback to refine action selection and output validity across iterations, enhancing the agent performance through linguistic feedback without need to update the weights of models.
In parallel, the scientist agent can coordinate multiple independent task-expert agents as tools, each designed as a specialized ReAct agent for a specific task or data source. These expert agents serve as domain specialists responsible for well-defined atomic tasks such as literature retrieval, online search, mathematical derivation, or data extraction. This functional decomposition ensures that each expert executes its task with minimal context switching, while the scientist maintains global task coherence and quality control.
III-B Multi-Agent Orchestration
The EpidemIQs framework is comprised of 5 phases. It initiates upon receiving a user query, entering the Discovery phase. In this phase, the DiscoveryScientist agent performs planning and reasoning to collect data relevant to the query following the Multi-Hop Question Answering (MHQA) paradigm [47]. Specifically, the data collection process is conducted in three iterative steps, with each step focusing on a distinct information source. This multi-step procedure enables the system to progressively refine its understanding, reduce errors, and enhance the overall accuracy of the gathered information[48]. The components of the Discovery team are as follows:
  • DiscoveryScientist coordinates the expert agents to collect relevant information, process the retrieved data, and synthesize the final output following a predefined structured format. This output encompasses a comprehensive description of the scenario, task specifications, and contextual information related to the disease and the affected population.
  • OnlineRetrieverExpert is responsible for acquiring relevant information from online sources. It generates and refines search queries in multiple stages, leveraging external APIs (e.g., Tavily API) to retrieve data from the web. The agent subsequently processes and integrates the retrieved information to generate a coherent response.
  • LiteratureRetrieverExpert conducts a multi-step search process using the Semantic Scholar API, based on the received query. For each retrieved query, the agent generates concise summaries and aggregates the most pertinent findings. The agent is prompted to generate the final summarized response based on multiple criteria, including relevance and accuracy.
  • MathExpert provides mathematical reasoning and analytical solutions for queries requiring quantitative analysis. The agent is capable of both generating and executing Python code, a capability demonstrated to enhance solution accuracy and mitigate hallucination risks ( [49] showed that it can increase up to 39.8% problem-solving capability over HumanEvals [50] benchmark). The backbone LLM of this agent should be an advanced language model with high reasoning capacity, able to perform chain-of-thoughts( such as OpenAI GPT o3-Jan-2025 or o3-mini-Jan-2025 models [51])
Following the completion of the Discovery phase, the Modeling phase is initiated to systematically construct the foundational components required for simulating epidemic dynamics on complex networks. This phase is orchestrated by a team of three specialized scientist agents, each responsible for a distinct modeling dimension: network topology, mechanistic model formulation, and parameterization. The functional responsibilities of these agents are delineated as follows:
  • NetworkScientist is tasked with generating a representative contact network that encapsulates the interaction patterns within the target population. The constructed network serves as a structural substrate upon which the epidemic process unfolds. Depending on the characteristics of the population and the epidemiological scenario, the network may adopt various configurations, including but not limited to regular networks or Erdős-Rényi random graphs [52], Barabási-Albert scale-free networks [53], and multilayer structures [54]. The NetworkScientist employs algorithmic procedures implemented in Python, utilizing scientific computing libraries such as NetworkX and SciPy, to model the network topology. The finalized network structure is preserved in the workflow repository for integration into the simulation pipeline.
  • ModelerScientist is responsible for defining the mechanistic framework of disease progression. It selects and configures an appropriate compartmental model based on the disease characteristics and insights derived during the Discovery phase.
  • ParameterScientist is invoked after the design of both network topology and compartmental model, to calculate and assign quantitative values to model parameters and initialize network states. It performs calculations through the Python coding tool.
Upon completion of the modeling phase, the framework transitions into the Simulation, wherein predictive experiments are conducted to analyze the modeled epidemic dynamics. This phase is executed by the SimulationScientist, which performs multiple stochastic simulations to capture the probabilistic evolution of contagion spread over the constructed contact network, and has delegated VisionExpert to verify successful outcomes of simulations. The VisionExpert has vision-language capabilities and is responsible for analyzing the visual outputs generated during the simulation phase. It interprets time-series plots or state evolution graphs, enabling interpretability and validation of dynamic behavior that may not be easily deduced from raw numerical outputs alone.
To conduct these simulations, the agent utilizes the Fast Generalized Epidemic Modeling Framework (FastGEMF) [55], a high-performance simulation engine capable of modeling arbitrary multi-compartment epidemic processes over multilayer network structures with computational efficiency. The SimulationScientist is pre-trained through one-shot learning examples to interface with FastGEMF. Throughout the simulation process, the agent records the generated time-series data using Python libraries such as Pandas and stores both the numerical outputs and corresponding state-evolution visualizations in the project repository.
Following the simulation stage, the framework advances to the Analysis, wherein a coordinated team of specialized agents derives actionable insights from the simulation outputs. This multi-agent analysis team is responsible for processing both numerical results and visual artifacts to extract context-specific metrics and interpret spread dynamics. The roles and responsibilities of the constituent agents are defined as follows:
  • DataScientist coordinates the specialized expert agents and integrates their outputs into a coherent interpretation of the simulation results. It identifies and computes key epidemiological metrics, such as peak infection time and size, epidemic duration, outbreak probability, etc., tailored to the nature of the modeled disease and the research query. The final output consists of numerical measures and qualitative analysis.
  • VisionExpert analyzes visualizations to extract metrics and interpret epidemic behavior (e.g., uncertainty bands around trajectories), that may not be easily interpretable from raw numerical data.
  • DataExpert is tasked with the interpretation of structured numerical data. This agent accesses simulation results stored in the system repository and applies analytical techniques using Python libraries such as Pandas, NumPy, and SciPy. Operating within a multi-hop paradigm, the DataExpert initially identifies the structure and content of the available data, then dynamically selects appropriate analysis strategies to compute requested metrics. This flexibility ensures the agent’s ability to generalize across varying data types and experimental settings.
The final phase of the framework is Automated Report Generation, wherein two coordinated expert agents synthesize the research findings into a structured academic manuscript. This process starts with the ReporterExpert, which is equipped with tools to access the outputs of the entire workflow, including simulation results, code scripts, and literature summaries. Guided by a predefined LaTeX-based manuscript template, the ReporterExpert sequentially drafts all major sections of the paper, such as the Title, Abstract, Introduction, Methods, Results, and Discussion. Following the generation of each section, the LatexCraftExpert is tasked to perform a syntactic and structural validation of the LaTeX code, resolving potential compilation errors and ensuring stylistic consistency with publication guidelines. Upon completion of the main draft, the LiteratureReviewExpert conducts a contextual relevance check, retrieving and integrating recent scholarly works aligned with the study’s findings. This agent augments the manuscript with a Background and Related Work section to strengthen the research academic positioning. Once the initial draft is completed, an iterative review loop is repeated by the ReporterExpert. It critiques each section to detect inconsistencies, redundancies, or logical gaps according to a user-defined number of revision cycles, $n_{max}$. Finally, it is compiled as a PDF file.
III-C Memory
Two specialized memory structures are considered for the agents. The scientist agent’s memory consists of short-term and long-term memory. The short-term memory encompasses all the current conversation details, including role description, input query, agent’s internally generated responses, interaction with tools, conversations with other agents, and the final output. The long-term memory stores all previous conversation history in a database, enabling semantic retrieval based on the current query.
III-D Structured Outputs
Typically, LLMs generate free-format texts as a natural language output, which does not follow a strict structure. However, the performance of the collaborative LLM agents requires precise and careful interoperability, as information sharing and communication can significantly impact overall outcomes [44]. Therefore, to make the data transition between each phase of the process robust, reliable, and more predictable, there are constraints defined for each team output structure, forcing them to produce output in a predefined structure (e.g., JSON schema). This not only facilitates data retrieval and improves prompt development efficiency, but it also enables automatic validation of each team’s output [56], ensuring all required fields are generated accordingly, and there is no missing data. If not, the agent retries until it fulfills the task or reaches the maximum number of tries allowed. While the output structure of each scientist contains different fields, all scientists share one common element: the reasoning and logical justification behind their choices and results. Each scientist is expected to defend their decisions against hypothesized criticisms, demonstrating why these were the most appropriate outcomes.
III-E Autonomous and Copilot Operation Modes
The EpidemIQs can perform in two modes: end-to-end autonomous and cooperative modes. The latter is named as a copilot, which can operate as an assistant to the user. The autonomous mode does not require human intervention, but rather initially provides the query. The copilot mode enables the human to intervene and provides feedback and instruction on the agent’s output, or asks for more actions to guide the agent toward the desired direction. After each phase, the human is asked to review the results, and they can either mark them as complete and proceed to the next phase or ask for new actions by providing comments.
III-F Implementation
The framework is implemented in Python using the Pydantic AI framework. Autonomous data acquisition is supported through API integrations using the Semantic Scholar API for literature retrieval and the Tavily API for web-scale information access. For multimodal processing, a Vision-Language Model (VLM) is implemented via the OpenAI SDK with Base64 decoding to enable visual data interpretation. Short-term memory utilizes the LLM’s context window, while long-term memory is stored in JSON and managed through Pydantic AI memory functions. Finally, we used Pydantic data models for schema validation and structural consistency across agent interactions.
IV Evaluation Approach
To assess the performance of the framework, we designed five questions, shown in Figure 4, with increasing depth and complexity.
Figure 4: Evaluation questions designed to assess the capabilities and limitations of the proposed framework across varying levels of complexity
The first question is the simplest, requiring an analytical investigation and comparison of stochastic simulations on heterogeneous versus homogeneous networks to assess how structural differences shape disease spread. The second focuses on identifying why the chain of transmission breaks and validating hypotheses through stochastic modeling. The third resembles the first but introduces a temporal network, which exceeds the scope of the provided simulation tools, testing system adaptability.
The fourth demands both analytical reasoning (or knowledge from literature) and an extension of the SIS model to a competitive $SI_{1}I_{2}S$ form on a multiplex network, exploring conditions for coexistence. The fifth is the most challenging, requiring analytical evaluation of random versus targeted vaccination strategies. Here, the key challenge is to notice the implicit constraint to ensure that the contact network contains enough degree-10 nodes to make targeted vaccination feasible while preserving its structural properties.
Collectively, these questions span the key aspects of network epidemics, such as topology, temporal structure, multilayer interactions and competing pathogens, termination mechanisms, network-aware interventions, physics-aware contact network, etc, providing a holistic evaluation framework. Only the first two questions were involved in the test and trials through the design of the EpidemIQs orchestration, and the last three were unknown to the framework.
Epidemiological research is inherently stochastic and complex, and admits multiple valid modeling approaches, leading to the absence of a unified standardized benchmark for evaluating the end-to-end epidemic research pipeline. Therefore, to evaluate our framework, we adopted an assessment strategy aligned with how the scientific community assesses research manuscripts through the peer review process. Consequently, we considered:
  • Human evaluation by peer review of written outputs and appraisal of deliverables. Specifically, we asked five blinded domain experts to review the manuscripts generated by EpidemIQs, based on the criteria in Supplementary Table S7.
  • LLM-as-Judge evaluation using GPT-4o to evaluate generated papers autonomously based on the same assessment criteria as human reviewers to deliver timely evaluation.
  • Completion success rate defined as the proportion of trials in which EpidemIQs autonomously executed the full workflow and produced a scientific report containing all pre-specified sections, independent of content quality.
  • Computational costs in terms of turnaround time, number of generated tokens, and total API cost.
Together, these measures capture both the scientific quality of the outputs and the reliability of the autonomous workflow. We compared our framework with a multi-turn single agent, such that an agent plans its next steps autonomously for up to the maximum number of steps, $s_{max}$. The single agent is instructed to perform all five phases with clear instructions (same as the agents in the EpidemIQs) and the same tools as in the multi-agent framework, along with VisionExpert to enable the agent to analyze the figures. This single agent needs to have an LLM with a long context window; therefore, we used OpenAI gpt-4.1-2025-04-14, as it has reasoning capability and support of structured output with a context window of about one million tokens, along with o3-2025-04014 as the current highly capable OpenAI model with a 200K tokens context window.
To prevent the prompt drift, in each turn we re-inject the main task into memory as **Reminder** main query is {query}. Do NOT lose sight of this, ensuring that the agent’s working context always contains the overarching objective. We considered the single agent as a baseline, and compared it with EpidemIQs’ implementation based on computational statistics, performance analysis, and review scores of the generated papers.
The setup used for assessment is presented in TableI.
TABLE I: Configuration Details for EpidemIQs
Parameter scientists experts MathExpert
LLM gpt-4.1- 2025-04-14 gpt-4.1-mini 2025-04-14 o3-mini 2025-01-31
$r_{\max}$ 1 - -
Retries 5 5 5
Tool Retries 50 50 50
Output Type Pydantic dataclass free-form free-form
V Results and Discussion
Across the five epidemic questions, EpidemIQs completed every phase of the pipeline as Discovery, Modeling, Simulation, Analysis, and Report writing in autonomous mode, achieving a 100% completion success rate in generating the final report file. Automated rubric scoring by the LLM‑as‑Judge assigned a mean quality score of 9.04±0.21 out of 10, while four network science Ph.D. students and one faculty reviewer gave an average expert rating of 7.98±0.35, praising the reports’ methodological soundness, clarity, and depth. Below, we summarize how the framework navigated each question, highlighting key decisions and results. The generated papers and a more comprehensive analysis are provided in the supplementary material.
Figure 5: Examples of paper generated by EpidemIQs for the second question.
Figure 6: An example of reasoning information NetworkScientist generated for question one
Question 1. In the first phase, analytical percolation theory and the literature confirm that outbreak behavior can differ according to the degree of heterogeneity. They suggest that, for a heterogeneous network, a scale-free graph, whose node-degree distribution follows a power law, producing a few highly connected hubs among many sparsely linked nodes, is an appropriate candidate to compare with a well-mixed homogeneous network, in which every node is connected to every other node, resulting in a fully connected graph. Because the question does not specify the intrinsic reproduction number, $R_{0}$, a value of 2.5 is adopted to reflect a disease that can be characterized by an SEIR model, such as COVID-19, for which $2<R_{0}<3$. Consequently, the deterministic model parameters are chosen to reflect typical values observed in COVID-19, incubation rate as $\sigma_{E\rightarrow I}=0.2/day$ (average incubation period of 5 days), recovery rate, $\gamma_{I\rightarrow R}=0.1/day$ (average infectious period of 10 days), and transmission rate $\beta_{S\rightarrow E}=0.25$. The framework then generates both graphs, initializes them with the five infected nodes, and tunes the transmission rate of the SEIR model on the basis of the intrinsic $R_{0}$ and a mean-field approximation to consider the effect of different network topologies.
In the simulation phase, three scenarios are executed for 300 stochastic realizations each: (i) random infectious seeding in the homogeneous network, (ii) random infectious seeding in the scale-free network, and (iii) infectious seeding at the highest-degree hubs in the scale-free network. Analysis of the simulation results corroborates the analytical theory. The fully connected homogeneous network exhibits a rapid, high-magnitude outbreak with near-complete infection penetration. By contrast, the scale-free network shows markedly different epidemic behaviour in both seeding schemes, with a much later epidemic peak, a lower peak prevalence, and a substantially smaller final epidemic size ($\approx 30\%$). Notably, hub-based infection seeding accelerates epidemic progression and yields a higher average final epidemic size, confirming the role of superspreaders in triggering more rapid outbreaks in networks with heavy-tailed degree distributions. Hence, the stochastic results confirm that degree heterogeneity decisively reshapes outbreak timing and magnitude. Complete description and results for this question are provided at Supplementary Section S1.1.
Figure 6 illustrates NetworkScientist’s reasoning and logic, exemplifying how each scientist justifies its task completion and the rationale behind its choices, thereby enhancing the interpretability of the outcomes.
Figure 7: An overview of the workflow for question four, highlighting each step’s goals, main actions, and outputs. The results have been truncated for clearer visualization. At the bottom, plots generated during the modeling and simulation phases are provided.
Question 2. Starting from the discovery phase, analytical results showed that the chain of transmission breaks when the effective reproduction number drops below unity, $R_{e}<1$. This can occur either when the susceptible population falls beneath a critical threshold or when the transmissibility of infectives decreases. Moreover, contact patterns (e.g., network topology) influence whether an outbreak sputters or ignites, thereby shaping its overall dynamics. To validate these findings, consistent with the surveyed literature, the framework employed an SIR model with two basic reproduction numbers: a supercritical case, $R_{0}=3$, and a subcritical case, $R_{0}=0.5$.
To isolate structural effects, two networks with identical average degree $\langle k\rangle=10$ and N=1, [^0] N=1{,}000 nodes were generated. The first is an Erdős–Rényi (ER) network representing homogeneous mixing, where each pair of nodes connects with probability $p$. The second is a scale-free Barabási–Albert (BA) network capturing heterogeneity; after an initial seed, each incoming node forms $m$ links to existing nodes, producing a heavy-tailed degree distribution.
Simulations began with ten randomly seeded infectives. For each network and $R_{0}$ value, transmission rates were adjusted (via mean-field approximation) to account for degree heterogeneity, and 75 stochastic realizations of each scenario were executed. In the supercritical scenario ($R_{0}=3$), the epidemic peaked rapidly and faded as susceptibles were depleted. In the ER network, the average susceptible count fell from 990 to $\approx 115$ and the recovered count rose to $\approx 885$, indicating herd-immunity termination. The BA network also exhibited fade-out but progressed faster and left a larger susceptible pool (final average epidemic size $\approx 520$, remaining susceptibles $\approx 480$), reflecting heterogeneity-mediated herd immunity.
For the subcritical scenario ($R_{0}=0.5$), intrinsic transmission inefficiency halted spread in both network topologies. Overall, the simulations confirm the analytical prediction that complete exhaustion of susceptibles is unnecessary; herd immunity can emerge when the susceptible fraction drops below a critical level or when per-contact transmissibility is sufficiently low. Figure 5 shows an overview of the generated paper, and details of the framework’s workflow are provided in the Supplementary Section S1.2.
Question 3. In the discovery phase, statistical analysis using a mean-field approximation derived expressions for the epidemic threshold in an activity-driven network, where each node activates at rate $\alpha=0.1$ and, upon activation, instantaneously forms $m$ edges, which are then removed in the next step. Because neither $m$ nor the recovery rate was specified, the experiment adopted a unit recovery rate and set $m=5$ as a plausible value. The analytical results indicated that, for $R_{0}=3$, a final epidemic size of 94% should be expected under these parameters.
To compare the analytical solution for the temporal network having 1000 nodes with its time-aggregated static counterpart, a static network was created by aggregating all temporal contacts over $T=1000$ time steps; for every node pair $(i,j)$, the edge weight equaled the total number of contacts during $T$. The recovery rate of one was chosen, and the infection rate was then calibrated for each network-driven SIR simulation so that they matched the mean-field derivations and were consistent with the respective network context (temporal or static). One random infectious node was selected as the initial condition, with all other nodes susceptible.
One hundred stochastic realizations were performed for each scenario, using FastGEMF for the time-aggregated static network and a custom discrete-time Markov-chain engine for the temporal network. The results revealed sharp contrasts in epidemic size and speed between the temporal and static scenarios. The temporal activity-driven network produced significantly smaller and slower epidemics relative to both the static-network simulation and the analytic predictions based on aggregated parameters. Specifically, the temporal SIR process yielded a mean final epidemic size of approximately 20.5% with a large standard deviation of 28.9%, indicating high stochastic variability and frequent epidemic die-out events at the chosen parameters. Such early stochastic die-out is ignored in mean-field ODEs, which assume deterministic branching of infections and never allow the process to fizzle when the first few seeds recover before transmitting the pathogen.
By contrast, the static-network simulation produced near-complete infection of the population (mean final size $\approx 81\%$) in a rapid fashion, closely approaching the analytical prediction of 94.05% derived from the classic mean-field equation. These results highlight the profound mechanistic effect of temporal causality on disease-transmission pathways. In the temporal network, edges represent instantaneous contacts that exist only transiently and are reorganized stochastically at each time step. Consequently, the temporal ordering of contacts restricts the accessibility of transmission chains, effectively increasing the epidemic threshold and reducing the reachable susceptible population. Many potential transmission paths present in the aggregated static network are simply infeasible in the temporal network because the required contact timings do not overlap to support causally coherent infections. Supplementary Section S1.3. provides the detailed outcomes for this question.
Figure 8: An overview of the workflow for question five, outlining each step’s goals, main actions, and outputs. For clarity, the results have been truncated. At the bottom, plots display the degree distribution, highlighting the large number of nodes with degree 10 and confirming that enough nodes are available for targeted vaccination, together with simulation results for four cases: no vaccination (baseline), vaccination below the critical threshold, at the threshold, and above the threshold.
Figures 7 and 8 display the results and process pipelines for questions 4 and 5, respectively. For the question 4 scenario, the framework could successfully find the criteria for the coexistence of the competitive exclusive viruses over multilayer networks, extend the analytical mean-field point of view that exists in literature, and show the dominance, coexistence, and extinction regimes in stochastic processes by designing the network properties that allow coexistence and tuning the parameters. Finally, for the last question, it not only determined the vaccination threshold for random and targeted cases but also carefully modeled the contact network to maintain the aforementioned centralities while having enough nodes to verify the targeted vaccination threshold. The presented results are truncated, and please refer to Supplementary Sections S1.4 and S1.5. for full detailed results of questions 4 and 5, respectively.
V-A Review Score and Computational Cost
TABLE II: Papers Evaluation of Five Epidemic Questions by AI and Human Reviewers
Question Topics
Readability
Relevance
& Motivation
Technical
Soundness
Experimental
Rigor
Limitation
& Ethics
Overall
AI Human AI Human AI Human AI Human AI Human AI Human
1. Effect of degree heterogeneity in SEIR model 9.5 7 9.5 9 9.5 8 10 8 9.5 8 9.5 8.0
2. Reasons of transmission break 9.5 7 10 10 10 6 9 10 8 9 9.7 8.2
3. Effect of temporality on epidemic behavior in activity-driven networks 9.5 9 9.5 9 9 7 9.5 7 8 7 9.2 7.8
4. Coexistence of exclusive bi-virus spread on multiplex networks 9 7 9 9 9.5 9 9.3 8 8.2 7 9.0 8.0
5. Herd-immunity through random and targeted vaccination in networks 9 6.5 9 9 10 8 10 9 8 8 9.2 8.1
Table II presents the evaluation scores assigned by an LLM-as-Judge and by human experts across five rubric criteria. For each of the five epidemic questions, the framework autonomously generated ten full papers, yielding 50 manuscripts. We randomly selected five of these for blind human review. Each reviewer scored every criterion on a 1‑to‑10 scale. Human scores averaged 7.98 with a standard deviation of 0.20 ($\approx 2.5\%$), indicating consistently strong performance by EpidemIQs. Reviewers praised the manuscripts’ comprehensive problem framing, robust experimental design and analysis, and balanced coverage of both analytical and stochastic aspects. The main criticisms concerned verbosity with occasional repetition of content, the use of uncommon vocabulary that slowed reading, and some missing or incomplete figure references.
Table III reports the average turnaround time, total number of tokens, and total cost (calculated by OpenAI’s May 2025 pricing) for all questions, measured from the initial query to the final report. A total cost of only \$1.57, a turnaround of under 30 minutes, a 100% success rate, and consistently high review scores demonstrate the strength of the framework and the efficiency of its multi‑agent orchestration. The low cost and rapid execution are achieved by splitting tasks between expert and scientist agents. Token‑heavy, low‑complexity jobs are delegated to fast, inexpensive models such as GPT‑4.1 mini, while more demanding works such as planning, reasoning, tool coordination, and output refinement is handled by the full GPT‑4.1 model.
TABLE III: Average Computation Statistics
Phase Time(sec.) Tokens(K) Cost(\$)
Discovery 158 98 0.1
Modeling 140 175 0.72
Simulation 198 88 0.23
Analysis 177 93 0.16
Report Writing 517 416 0.36
Overall 1190 870 1.57
V-B Multi-Turn Single-Agent vs EpidemIQs Orchestration
In this section, we ran the single-agent across the same five questions ten times, using two OpenAI LLMs, GPT‑4.1 and o3. Single-agent‑GPT‑4.1 achieved an average completion success rate of 78±7.7% and an average human-review score of 5.06, whereas single-agent-o3 reached 80.0±6.32% with a human-review score of 5.68. The LLM-as-Judge provides comparative evaluations of the successful papers generated by single-agent and EpidemIQs. Table IV presents these scores for all questions.
TABLE IV: Overall Score for AI and Human Expert Evaluation of Generated Papers
Model question 1 question 2 question 3 question 4 question 5
AI Human AI Human AI Human AI Human AI Human
EpidemIQs 9.5 8.0 9.5 8.2 9.2 7.8 9.0 8.0 9.2 8.1
single-agent-GPT4.1 8.9 5.4 7.3 5.3 8.9 4.8 7.5 4.6 8.4 5.2
single-agent-o3 8.5 5.8 8.6 6.3 8.3 5.5 8.6 5.8 7.4 5.2
Surprisingly, the AI-generated scores differ only slightly across the three cases, while the human reviews reveal substantial disparities in their outcomes. The single-agent-GPT‑4.1 setup addressed the first three questions to some extent; however, it was limited: it handled the analytical component inadequately, ran only a few stochastic simulations, and in some instances merely solved the ODE system instead of performing exact stochastic simulations. For the more complex question 5, none of the trials identified the target vaccination threshold or model the network appropriately. Furthermore, the parameters and the rationale for their selection were not discussed comprehensively.
The conclusions were sometimes misleading, as they were not based on the simulation results; instead, the agent often relied on prior knowledge or literature. When grounded in results, they were still occasionally inaccurate, for example, in the targeted-vaccination case, the agent typically claimed that vaccinating all high-degree nodes was necessary, largely because it generated the network before calculating the analytical thresholds, resulting in very sparse high-degree nodes. Human reviews also mention that while the agent could address the task to some degree, it is in shallow depth, missing figures, insufficient simulations, and limited analyses, and the overall failure to address the tasks comprehensively, compared to rigorous analysis and experimentation in the MA setup.
On the other hand, The single-agent-o3 did better job with analytical aspect of the question and better writing quality, however, the simulations still was very limited, and the agent shows more autonomy in how to accomplish tasks rather based on the expert knowledge provided as instructions, for example for stochastic simulation it did not always use the engine provided for it or ignoring literature and online searches. Though the single-agent-o3 generally generated higher quality reports, there was still some cases when there were inaccuracies, for example for question four, even the coexistence was not achieved through stochastic simulation, it was concluded that simulation results confirm that coexistence is possible when analytical conditions are met or the agent mentioned the simulation were performed by FastGEMF over 1000 realizations, however scripts in the repository showed custom engine for deterministic simulation was used to simulated SI instead of SIR, which showed a huge drift from the main task and hallucination of the agent in the report. Also, in none of the cases could it correctly address the targeted vaccination in question five. Finally, while reviewers praised the writing quality, conciseness, and analytical aspects of the reports, they found that they fell short in having enough depth, simulation rigor, and failure to address complex tasks comprehensively.
It is also necessary to mention that, to avoid violating OpenAI’s usage policy when using the o3 model, we changed the questions from virus spread to meme spread. Otherwise, the success rate of single-agent-o3 was less than 30%. We did not consider this API error in our benchmarks for evaluating the o3 success rate. It is important to emphasize that this change in context does not affect the validity of the comparative results, as epidemic spread can easily be generalized for social cases such when a meme spread over a population.
Overall, the human reviews and performance analysis demonstrated the superior performance of EpidemIQs compared to the single-agent implementation. Single-agent systems consistently failed to address complex scenarios, exhibited drift in their outcomes, lacked depth in their reports, and had a significantly lower success rate in completing tasks. Table V presents the computational statistics for these three different implementations. These comparisons are particularly meaningful given that the EpidemIQs achieved its superiority at a cost comparable to single-agent-GPT4.1 and at less than half the cost of single-agent-o3, despite the o3 model having greater reasoning and coding capabilities than those employed in the EpidemIQs, as mentioned in Table I. Moreover, the superiority of the proposed framework extends beyond performance and computational efficiency, offering greater interpretability by providing explicit reasoning for each step.
TABLE V: Comparison of Average Computation Statistics Across All Questions Over Three Different Implementations
Method Time (s) Tokens (K) Cost (\$)
EpidemIQs 1190 870 1.57
single-agent-GPT4.1 214 312 0.91
single-agent-o3 436 165 4.13
V-C Discussion, Limitations, and Ethical Concerns
Although the EpidemIQs framework consistently produces high‑quality manuscripts throughout the workflow, it is not designed to replace human authorship in scientific research, but rather to be an assistant to human experts to facilitate the research process for testing and implementing ideas in the epidemic modeling field. EpidemIQs can still make mistakes, and AI‑generated prose can be highly persuasive, as [57, 58, 59] discussed, so human oversight remains essential to safeguard accuracy and integrity, and generated papers are meant to provide a comprehensive and unified result of the framework, making it easy to analyze the outcomes.
Across all of our experiments, we found no major instances of hallucination in EpidemIQs outputs, and performance was remarkably consistent. Since LLMs are non‑deterministic, they may tackle the same question slightly differently each time; however, no substantial drift or incorrect reasoning was observed. For example, in Question 5, ModelScientist sometimes designs an SIRV (V represents vaccinated state) model and at other times an SIR model, or NetworkScientist may construct networks with different degree distributions yet the same target centrality. These variations are valid and do not diminish performance; they simply reflect the natural diversity of approaches to a single problem. However, there were errors observed during the Analysis phase. For example, in question three, where DataExpert misinterpreted an average recovery count below one as a fraction rather than a discrete number of individuals, DataScientist detected this inconsistency and deferred to VisionExpert, whose interpretation matched both the simulation protocol and analytical expectations. While such checks limit the impact of isolated calculation errors, occasional mistakes by DataExpert indicate that automated data analysis may be a part of further improvement.
The framework also demonstrated sound awareness of its computational constraints. For example, in question three, we explicitly restricted simulations to the FastGEMF engine. SimulationScientist adhered to this limitation, employing FastGEMF exclusively for static networks, and it was explicitly noted in the final report that software limitations prevented temporal network experiments. This transparency is critical, as it prevents the framework from falsely claiming task completion. In contrast, such issues were observed in the single-agent implementation, where it falsely hallucinates task accomplishment. Furthermore, the reliability of scientific outcomes often hinges on the algorithms employed, which can be inherently complex. For instance, in one of the trails within question three, SimulationScientist designed a custom engine for simulating disease spread over a temporal network but erred in updating the order of states. This led to the premature recovery of initially infected nodes, preventing further spread and resulting in an underestimation of the final outbreak size and probability. Although DataScientist noted a significant discrepancy between simulated and expected analytical results, it incorrectly attributed the difference to the effects of temporality rather than the engine’s design flaw.
To further investigate the agents’ attention to scientific details, in question three we explicitly specified the constant activity rate (and not activity probability per unit time), $\alpha=3$, which is unlikely in real-world case studies and rare in literature but was chosen to test the agents’ attention to subtle details. However, NetworkScientist creates the network edges with the probability of a node being activated as $p=\alpha\Delta t$, with $\Delta t=1$, as the first order Taylor series approximation of $p=1-\exp^{-\alpha\Delta t}$, where is only is acceptable if $\alpha\Delta t\ll 1$ (However, if $\alpha$ is probability $\alpha\Delta t$ is exact), making it invalid for $\alpha\Delta t=3>1$ resulting all nodes being activated each time, instead of having exact probability $\approx 0.95$.
These observations emphasize two points:
1) the importance and impact of providing appropriate tools to obtain consistent and reliable results, and 2) the continued need for a human expert to validate results, because agents can still make mistakes that are difficult to detect and may produce persuasive responses that justify incorrect conclusions.
Our framework is unique, with no direct comparators. However, a comparison with The AI Scientist, a general tool for scientific discovery from ideation to paper writing [34], for question one, showed the failure of The AI Scientist. It analyzed the ODE-based analytical component but struggled with stochastic simulations over the network, producing hallucinated results and conclusions. It is essential to mention The AI Scientist performance was mostly evaluated on AI-related subjects, such as diffusion models or transformers, emphasizing that this comparison does not imply overall superiority of one framework over another.
Summarizing the main limitations of the current work, we have:
  • Although no major hallucinations were found in the results, a few hallucinated references did appear in the final report. Moreover, while the reviewers praised the papers’ comprehensiveness, they also noted some repetitive content that lengthened the reading time, and sometimes tables were so large that portions extended beyond the page, suggesting enhancing the report generation part.
  • The performance of the EpidemIQs is highly dependent on the LLM models, while our framework is model agnostic, as long as these models can support structured outputs and tool calling, the results can vary based on the LLM choices.
  • The literature review relies only on the abstract and summary of key findings, rather than delving deeply into the full content of the papers. Future frameworks capable of thoroughly analyzing the full material of each paper could enhance and benefit from deeper data retrieval.
  • The current framework focuses on epidemic modeling over complex networks as a proof of concept. Enhancing it to broader methods such as Agent-Based Models(ABM)[60, 61], individual-based [60], or statistical and data-driven approaches [60, 62] can significantly increase the applicability of the framework.
  • The agents’ performance hinges on the tools at their disposal. As noted in question three, their performance and reliability decline when they encounter highly technical problems without the requisite expert knowledge and appropriate tools. This shortcoming explains why the general‑purpose module proposed by [34] fails in question one.
  • AI evaluation in the automated review process shows divergence from human expert reviews, aligning with other works’ findings, such as Agent Laboratory [30]. However, contrasting with near-human performance reported in [34], it suggests that human involvement is necessary. While an LLM-as-Judge can be beneficial for providing high-level insights, it cannot be fully trusted, since it relies on superficial patterns rather than robust analysis criteria.
  • We did not include ideation in our framework, primarily because epidemic research is highly complex and interdisciplinary. Generating new research ideas requires careful consideration of resources, advanced tools (see Limitation 4), expert knowledge across multiple domains, and access to diverse data sources, such as mobility, weather, and epidemiological data. For example, in vector-borne diseases, the populations of mosquitoes or birds play a critical role in epidemiological modeling. To the authors’ knowledge, such ideas can quickly become too complex for current agentic AI systems to handle appropriately, given concrete bottlenecks such as the lack of structured access to high-quality mobility/weather/entomological data, difficulty in reasoning across coupled mechanistic models, or access to specific tools. Therefore, we focused on specific network-based epidemic modeling tasks as a proof of concept to showcase the framework’s potential benefits.
Therefore, although EpidemIQs offers great potential to accelerate epidemic-modeling research and allows researchers to focus on high-level ideation, this promise also introduces new ethical challenges that must be taken seriously. Its ability to autonomously generate high-quality papers, supported by analytical insights, rigorous simulations, and comprehensive analyses, may inadvertently lower the barriers to producing low-standard scientific work, potentially overwhelming the peer-review system and jeopardizing the integrity of academic discourse. Moreover, as noted earlier, the automated process may introduce mistakes in algorithm design that are implicit and difficult to detect, or it may yield results biased by its underlying knowledge or algorithms, leading to skewed outcomes. These concerns underscore the necessity of disclosing AI involvement in the research process to mitigate such risks and maintain accountability, fairness, and transparency.
VI Data Availability
All the data and results discussed in this paper are provided as supplementary materials. The EpidemIQs is also provided as an open-source Python module at github.com/will-be-announced-upon-publication.
VII Conclusion
In this study, we present EpidemIQs, a multi‑LLM agent framework that harnesses the reasoning, planning, and problem‑solving abilities of LLMs and integrates them with online and literature retrieval, multimodal data processing, code generation and execution, stochastic simulation, and visual and data analysis. These characteristics enable it to autonomously carry out every key stage of epidemic modeling research, from discovery through modeling and simulation to analysis and, finally, the synthesis of findings into a full scientific report. This entire process takes less than half an hour, whereas it would normally take an average graduate-level student days to accomplish the same task.
The evaluation of results across questions addressing various aspects of network-based epidemic modeling demonstrates that the framework consistently maintains high performance across our trials. Consequently, EpidemIQs can extend existing analytical approaches to stochastic processes, infer implicit constraints in research questions, and design scenarios that address cases beyond the immediate scope of available tools and expert knowledge.
The comparative analysis of EpidemIQs and the single-agent implementation consistently showed that EpidemIQs outperformed the single-agent variants across all tasks. It achieved an 100% success rate average human‑evaluation score of 7.98 at a cost of \$1.57, whereas the single-agent configurations based on GPT‑4.1 and o3 obtained scores of 5.06 and 5.68 at costs of \$0.91 and \$4.13, respectively.
At present, EpidemIQs focuses exclusively on epidemic modeling over networks as a proof of concept; nevertheless, it has already demonstrated the potential of deploying LLM agents in a collaborative environment. This opens an exciting path for further performance gains by integrating advanced tools, such as deep learning techniques, probabilistic methods, and data-centric models, heralding a promising future for epidemic modeling research. Moreover, because the framework’s accuracy and performance depend heavily on the underlying LLMs, the rapid advances in this field are likely to yield substantial improvements for the entire multi agent system, as our framework is designed to be model agnostic.
EpidemIQs, while demonstrating strong performance and producing high‑quality output, is not intended to, and should not, replace human authorship in the scientific process, as it still makes mistakes and should not be fully trusted; it is best regarded as a highly capable assistant. Researchers in epidemic modeling can use it to test and implement their ideas quickly and at low cost, freeing them from time‑consuming tasks such as setting up simulations or writing repetitive code and allowing them to focus on the conceptual and creative aspects of their work.
Future work should consider the broader applications of epidemic modeling by integrating real-world data and other advanced computational tools into the model design to address open-ended real-world problems and evaluate its applicability for forecasting current outbreaks. Additionally, they should address ethical concerns regarding the misuse or dual use of such a framework, as automated modeling and report generation could be exploited to produce misleading forecasts or scientifically plausible misinformation, intentionally or unintentionally influencing public perception or policy decisions. The system’s ability to simulate outbreak dynamics at scale may also lower barriers for malicious actors to explore harmful scenarios or identify vulnerabilities in public health defenses. Additionally, integration with sensitive epidemiological data poses privacy risks, emphasizing the need for implementing appropriate safeguards.